OpenStudy (anonymous):

can some one teach and help me understand how to slove this better?? two garden plots are to have the same area. one is square one is retangular. the rectangular plot is 4 meters wide and 16 meters long. how large is one side of the square garden plot in meters?

OpenStudy (anonymous):

Okay, first, we have two gardens  Let's make the sides of the square equal to "a" and the sides of the rectangle equal to "c" and "d" now, the area of the square is given by \[A_1=a*a\] and the area of the rectangle \[A_2=c*d\] And we know that \[A_1=A_2\] if c and d are equal to 4 and 16 \[A_1=a*a=4*16\] that means that \[A_1=a^2=64\] now solve for a and you will know the length of a side of the square c:

OpenStudy (anonymous):

would i divide or multiply ???

OpenStudy (anonymous):

how do you solve x^2=4 ??

OpenStudy (anonymous):

divide cuz x=2 if u divide 4 by 2 ud get 2 then 2*2 wud equl 4

OpenStudy (texaschic101):

area of a rectangle is L x W so the area of this rectangle is 4 x 16 = 64 m area of a square is s^2 but you want one side, so the perimeter of a square is 4x (4 times each side) 4x = 64 x = 64/4 x = 18 m

OpenStudy (texaschic101):

wait...that might not be right...hold on

OpenStudy (anonymous):

nope, you aply a square root \[x^2=4\] \[\sqrt{x^2}=\sqrt{4}\] then \[|x|=2\] so x=2 or x=-2 In our case, \[a^2=64\] we apply the square root \[\sqrt{a^2}=\sqrt{64}\] \[|a|=8\] so a=8 or a=-8 BUT a is a lenth and there're not negative lengths so, a=8 Did you get it?

OpenStudy (anonymous):

so theres no dividing just squaring???

OpenStudy (texaschic101):

64 = x^2 so your answer is the square root of 64 = 8....Umangiasd is correct

OpenStudy (anonymous):

cuz a rectangle IS a square??

OpenStudy (texaschic101):

a rectangle is not a square

OpenStudy (anonymous):

look at what i draw, and think why I did what i did.

OpenStudy (texaschic101):

the area of a square is x^2.....and we know the area is 64 because it is the same as the area of the rectangle. So, 64 = x^2. To find x we then need to find the square root of 64, which is 8

OpenStudy (anonymous):

okaii so i did divide and both a's equal 8 cuz 8*8 equals 64

OpenStudy (anonymous):

i get it now :) thanks